Rose Berry is attempting to evaluate 2 possible portfolios, which consist of the same 5 assets held in different proportions. She is particularly interested in using beta to compare the risks of the portfolios, so she has gathered the data shown in the following table:

                                                                                            Portfolio Weights

Asset                     Asset beta                    Portfolio A              Portfolio B

1                               1.30                             10%                         30%

2                               .70                                30                           10

3                               1.25                              10                           20

4                               1.10                              10                           20

5                                .90                               40                           20

Totals                                                            100%                     100%

a. Calculate the betas for portfolio A and B

b. Compare the risks of these portfolios to the market as well as to each other. Which portfolio is more risky?                
Answer #1
Asset Beta Portfolio A Portfolio B
1 1.3 10% 30%
2 0.7 30% 10%
3 1.25 10% 20%
4 1.1 10% 20%
5 0.9 40% 20%
a) Beta A 0.935
    Beta B 1.11

Beta can be found out using the sumproduct function of excel:

Beta A = 1.3 * 10% + 0.7 * 30% + 1.25 * 10% + 1.1 * 10% + 0.9 * 40% = 0.935

Beta B = 1.3 * 30% + 0.7 * 10% + 1.25 * 20% + 1.1 * 20% + 0.9 * 20% = 1.11

b) Beta measures the systematic risk of the portfolio, that is, the senistivity of the portfolio to the market.

Since Beta of portfolio B is more than that of portfolio A, Portfolio B is more risky than Portfolio A.

Comparison with the market

Market Beta = 1

Beta A = 0.935

Beta B = 1.11

Since Beta of portfolio A is less than the market beta, portfolio A is less risky than the market portfolio.

Since Beta of portfolio B is more than the market beta, portfolio B is more risky than the market portfolio.

The order/ranking of portfolios in terms of riskiness measured using Beta is as follows:

1) Portfolio B

2) Market Portfolio

3) Portfolio A